eva invested a certain amount of money at 15% intrest and $2500 more than that amount at 16%. her total yearly intrest was $1485. How much did she invest at each rate?
Answer by stanbon(75887)

eva invested a certain amount of money at 15% interest and $2500 more than that amount at 16%. her total yearly interest was $1485. How much did she invest at each rate?
-------------------
Equation:
interest + interest = interest
0.15x + 0.16(x+2500) = 1485
----
Multiply thru by 100 to get:
15x + 16x + 16*2500 = 148500
----
31x = 108500
x = $3500 (amt. invested at 15%)
----
x+2500 = $6000 (amt. invested at 16%)
==========================================
Cheers,
Stan H.
================